OpenStudy (anonymous):

Hey I have this problem on Linear equations I can't figure out, anyone able to help?

OpenStudy (dan815):

ok but can I have a medallaion

OpenStudy (anonymous):

it would help if you posted the problem

OpenStudy (dan815):

is it linear differential equation or matrix linear equations

OpenStudy (dan815):

homogenous linear equations?

OpenStudy (lifeisadangerousgame):

Here's the file guys

OpenStudy (dan815):

omg GOOD BYE

OpenStudy (dan815):

just kidding!

OpenStudy (dan815):

basically the relationship between perpendicular lines is their slope first write equation in y=mx+b form and look at the m if m =a/c then the other m should be -c/a

OpenStudy (lifeisadangerousgame):

Ohh okay, I've got it! :D Thank you everyone!

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq3.htm

OpenStudy (dan815):

:)

OpenStudy (dan815):

ill give u an example to clarify just in case say y = 2x+b1 so m = 2 or otherwise written as 2/1 then a line perpendicular to it wud have the slope -1/2 so y = -1/2x + b2 wud be perpendicular

OpenStudy (lifeisadangerousgame):

That makes sense, thanks! ^_^ I'm gonna go work on some practice problems just to double check :b Thanks again for replying so quickly! :D

OpenStudy (dan815):

no come back to me when u wanna ask me a riddle question i love riddle questions :)

OpenStudy (lifeisadangerousgame):

Haha will do! :D

OpenStudy (anonymous):

Thank you guys so much this really helped!